OpenStudy (anonymous):

4(x+2)-(x+3)=8 Please explain to me

OpenStudy (nubeer):

first multiply 4 with numbers inside 4x +8 - x+3 = 8 3x +11 = 8 3x = -3 x =-1

OpenStudy (jiteshmeghwal9):

4(x+2) 4x+8 so, (4x+8)-(x+3)=8 4x+8-x-3=8 3x+5=8 3x+5-5=8-5 ....(transposition) 3x=3 x=3/3=1

OpenStudy (anonymous):

4(x+2)-(x+3)=8 4x + 8 - x - 3 = 8 4x - x = 8 - 8 + 3 3x = 3 x = 1

OpenStudy (anonymous):

Thanks all of you!

OpenStudy (jiteshmeghwal9):

yw :)

OpenStudy (anonymous):

You are welcome